OpenStudy (anonymous):

PLEASE CHECK ANSWER !A! MEDAL MEDAL MEDAL Solve, using linear combination. 5x + y = 2 4x + y = 4 (–3, 17) (–2, 12) (0, 2) (1, –8)

OpenStudy (nuke_ur_friend):

witch 1 is ur awnser

OpenStudy (nuke_ur_friend):

???

OpenStudy (anonymous):

A

OpenStudy (nuke_ur_friend):

oh ok then u r corccet i tink

OpenStudy (anonymous):

hm thats why i dont know if im correct either ?

OpenStudy (nuke_ur_friend):

hold up hold up hold up

OpenStudy (anonymous):

okaay

OpenStudy (nuke_ur_friend):

lol

OpenStudy (anonymous):

so would that be like substitution or elimination? if so... subtract the second equation from the first equation to get rid of y: 5x + y = 2 4x + y = 4 x=-2 (now sub that into with equation to get y) 5(-2) + y = 2 y=12 so the answer is B

OpenStudy (anonymous):

np

OpenStudy (nuke_ur_friend):

ya it is B.) not A.)

OpenStudy (anonymous):

*sub into either equation

OpenStudy (anonymous):

oh thaanks @Nuke_ur_friend but @rimshaa gave an answer first so thannnks to the both of you (:

OpenStudy (anonymous):

no prob

OpenStudy (nuke_ur_friend):

idk i just like 2 help :)

OpenStudy (anonymous):

thanks . would you stay and help me with another prob ?

OpenStudy (anonymous):

or close this and open another one so i can give you a medal to ?

OpenStudy (nuke_ur_friend):

ya i would love 2

OpenStudy (nuke_ur_friend):

help u

OpenStudy (nuke_ur_friend):

agin lol

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

How many solutions are there to the following system of equations? 2x + 5y = –11 10x + 25y = –55 (Points : 4) 0 1 2 infinitely many

OpenStudy (anonymous):

This answer i am hoping it would be B ? ?

OpenStudy (nuke_ur_friend):

it is 1 dats wut my bro told me i had a problem like dat befor

OpenStudy (anonymous):

Thanks (: !

OpenStudy (nuke_ur_friend):

lol yup

OpenStudy (anonymous):

2x + 6y = –12 10x + 32y = –62 0 2 1 infinitely many

OpenStudy (anonymous):

What grade are you in ?@Nuke_ur_friend

OpenStudy (anonymous):

& i dont have an answer for that one ? ? ?

OpenStudy (nuke_ur_friend):

well idk but i'm 16 lol i just go wit life lol

OpenStudy (nuke_ur_friend):

i think it is 0

OpenStudy (anonymous):

oh kuu kuu thanks for the help like foreal doe ! Maybe ill become a fan of your's since i couldnt give a medal (:

OpenStudy (nuke_ur_friend):

ok i would ♥ love dat ill fan u 2 k

OpenStudy (anonymous):

ok ok !

OpenStudy (anonymous):

maybe then again i can ask you for help ?

OpenStudy (nuke_ur_friend):

and do u like vollie ball or wut ever

OpenStudy (anonymous):

volleyball . yees . how you know ?

OpenStudy (nuke_ur_friend):

sorr sticke kys

OpenStudy (nuke_ur_friend):

littl sis spillted soda on it

OpenStudy (anonymous):

is their another way we can chat because ha people can see this . lol

OpenStudy (anonymous):

????

OpenStudy (nuke_ur_friend):

and i knew because u look like u like vollie vall

OpenStudy (nuke_ur_friend):

pm

OpenStudy (anonymous):

oh well thanks i love volleyball . & i dont know how to pm ? you pm me ?

OpenStudy (nuke_ur_friend):

or e mail me at corycaulkins13232@gmail.com my little bro e mail he gave me

OpenStudy (anonymous):

lolk will do .. when doe ?

OpenStudy (nuke_ur_friend):

lol ok ill be on till 12:00 or 1:00

OpenStudy (anonymous):

oh hahaha kk well iwill email you on gmail cause im going off this .. thaanks for the help

OpenStudy (nuke_ur_friend):

ook bye !!!!